# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ Appendix A Reused semantic resources
Table 16: Reused semantic resources, the systems that reuse them, and the challenges they address.
@@
Amigo Device Ontology Provides support for the description of devices, user context, quality of service parameters, and communication protocols. [92] Interoperability; Context awareness
Association Ontology [^39] http://purl.org/ontology/ao/core Provides basic concepts and properties for describing specific association statements to something. [110] Interoperability
Basic Formal Ontology [^40] https://basic-formal-ontology.org A foundational ontology designed for use in supporting information retrieval, analysis, and integration in scientific and other domains. [84] Interoperability
Basic Geo (WGS84 - lat/long) [^41] https://www.w3.org/2003/01/geo A basic RDF vocabulary that provides a namespace for representing lat(itude), long(itude), and other information about spatially-located things. [73] Interoperability; Context awareness
@@
Chronic Obstructive Pulmonary Disease Ontology (COPDology) Contains concepts related to the disease, environment, equipment, patient data, and treatment. [100] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support; Explainability
Cyc Ontology [^42] https://cyc.com/platform/ A large knowledge base of common sense and background knowledge. [104] Interoperability
Data Analytics for Health and Connected Care (DAHCC) Ontology Captures care, patient, daily life activity recognition and lifestyle domain knowledge. [83, 75] Interoperability; Context awareness; Situation detection; Situation prediction; Explainability
@@
Diabetes Mellitus Diagnosis Ontology A diabetes knowledge base that supports automatic reasoning for solving problems related to diabetes diagnosis. [84] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
@@
Diseasome A.k.a. the human disease network; a graph that captures all genetic disorders and disease genes and the links between them. [72] Interoperability; Situation detection; Situation prediction; Explainability
@@
FHIR & SSN-based T1 Diabetes Ontology (FASTO) Ontology integrating FHIR standard and SSN ontology with clinical practice guidelines for real time management of insulin for diabetes patients. [77] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
Food Ontology A lightweight ontology for describing recipes, ingredients, menus, and diets. [106] Interoperability; Decision support; Explainability
Friend of a Friend (FOAF) Ontology An ontology describing persons, their activities, and their relations to other people and objects. [85, 90, 100, 73, 106, 107, 110] Interoperability; Context awareness
Table continued on next page.
Reused semantic resources, the systems that reuse them, and the challenges they address.
General User Model Ontology A user model ontology for use in user-adaptive or ubiquitous computing systems. [100] Interoperability; Context awareness
@@
HealthIoT Ontology Provides semantic representation of both the medical connected objects (i.e. sensors) and their data. [94] Interoperability; Context awareness
@@
International Classification of Functioning, Disability and Health (ICF) A framework for describing and organising information on functioning and disability. [106] Interoperability; Situation detection; Explainability
@@
IoT-lite Ontology A lightweight ontology to represent IoT resources, entities, and services. [105] Interoperability; Context awareness
@@
Medical Subject Headings (MeSH) A vocabulary for indexing, cataloging, and searching biomedical and health-related information. [73] Interoperability
MIMU-Wear Ontology Describes wearable sensor platforms consisting of mainstream magnetic and inertial measurement units. [94] Interoperability; Context awareness
Mobile Object Ontology Models temporal, spatial, and domain related information about mobile objects. [105] Interoperability; Context awareness
@@
OwlSpeak Ontology Represents static and dynamic concepts related to spoken dialogue. [100] Interoperability; Context awareness
OWL-S Ontology Describes Semantic Web services for automatic service discovery, invocation, composition, and interoperation. [87, 92] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support
OWL-Time Ontology Describes the temporal properties of resources. [83, 100, 105, 107, 110] Interoperability; Context awareness
Parkinson and Movement Disorder Ontology Captures neurological findings, treatment plans, and instruments used to evaluate various traits of Parkinson’s Disorder. [75] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support; Explainability
Physical Activity Concept Ontology Captures various concepts that are required to describe one’s physical activity. [94] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support; Explainability
@@
Table continued on next page.
Reused semantic resources, the systems that reuse them, and the challenges they address.
@@
Semantic Sensor Network (SSN)/Sensor, Observation, Sample, and Actuator (SOSA) Ontology Describes sensors and their capabilities, measurement processes, observations, and deployments. [81, 83, 84, 85, 90, 94, 105, 74, 107, 75] Interoperability; Context awareness
@@
SWRL Temporal Ontology Defines defines a set of builtins that can be used to represent temporal information in SWRL rules. [84] Interoperability; Context awareness
Systematized Nomenclature of Medicine Clinical Terms (SNOMED CT) A collection of medical terms providing codes, terms, synonyms and definitions used in clinical documentation and reporting. [81, 84, 98, 99, 73, 105, 107, 71] Interoperability
@@ Appendix B Challenges assessment
# System Year Interoperability (score out of 6) Cont. awareness (score out of 7) Sit. detection (score out of 3) Sit. prediction (score out of 3) Dec. support (score out of 6) Explainability (score out of 6) Unc. handling (score out of 6) Overall (cumulative score out of 34)
1 Akhtar et al. [76] 2022 Low (1) High (5) Medium (2) ✗ (0) Medium (3) Medium (3) Low (1) Medium (15)
2 Ali et al. [77] 2021 Medium (3) Medium (4) Very High (3) ✗ (0) Low (2) Low (2) Low (1) Medium (15)
3 Ali et al. [78] 2020 Low (2) Medium (3) Medium (2) Low (1) Medium (3) Low (1) Medium (2) Medium (14)
4 Ali et al. [79] 2018 Low (1) Medium (3) Very High (3) ✗ (0) Medium (3) Low (2) Medium (2) Medium (14)
5 Alti et al. [80] 2022 Low (1) High (5) Medium (2) Low (1) Low (1) Low (2) ✗ (0) Low (12)
6 Chatterjee et al. [81] 2021 Medium (4) High (5) Medium (2) ✗ (0) Medium (2) Medium (3) ✗ (0) Medium (16)
7 Chiang and Liang [82] 2015 Low (1) High (5) Medium (2) Low (1) Medium (4) Medium (3) Low (1) Medium (17)
8 De Brouwer et al. [83] 2022 Medium (4) Medium (6) Very High (3) Low (1) Medium (3) Low (2) Low (1) Medium (20)
Table continued on next page.
Summary of the extent to which each system addresses the seven key challenges.
# System Year Interoperability (score out of 6) Cont. awareness (score out of 7) Sit. detection (score out of 3) Sit. prediction (score out of 3) Dec. support (score out of 6) Explainability (score out of 6) Unc. handling (score out of 6) Overall (cumulative score out of 34)
9 El-Sappagh et al. [84] 2019 High (5) High (5) Low (1) ✗ (0) Medium (3) Low (2) ✗ (0) Medium (16)
10 Elhadj et al. [85] 2021 High (5) High (5) Medium (2) ✗ (0) Low (2) Medium (4) ✗ (0) Medium (18)
11 Esposito et al. [86] 2018 Low (1) Medium (4) Very High (3) ✗ (0) Low (2) Low (2) Medium (2) Medium (14)
12 Fenza et al. [87] 2012 Low (2) Low (2) Very High (3) Low (1) ✗ (0) Low (1) Medium (2) Low (11)
13 Garcia-Valverde et al. [88] 2014 ✗ (0) Medium (3) Low (1) ✗ (0) Medium (2) Low (2) Low (1) Low (9)
14 Hadjadj and Halimi [89] 2021 Medium (3) High (5) Low (1) ✗ (0) Medium (3) Low (2) ✗ (0) Medium (14)
15 Henaien et al. [90] 2020 Medium (4) Medium (4) Low (1) ✗ (0) Medium (3) Low (2) ✗ (0) Medium (14)
16 Hooda and Rani [91] 2020 Low (1) Low (2) Medium (2) ✗ (0) Low (2) Low (2) Low (1) Low (10)
17 Hristoskova et al. [92] 2014 Medium (3) High (6) Medium (2) Low (1) Medium (3) Low (2) Low (1) Medium (18)
18 Hussain and Park [93] 2021 Low (2) Medium (3) Medium (2) Low (2) Medium (2) ✗ (0) Low (1) Low (12)
[…]
21 Khozouie et al. [96] 2018 Low (2) High (5) Low (1) ✗ (0) Low (1) Low (2) ✗ (0) Low (11)
22 Kim et al. [97] 2014 Low (1) Medium (3) Very High (3) ✗ (0) Low (1) Low (1) ✗ (0) Low (9)
23 Kordestani et al. [98] 2021 Medium (3) Medium (4) Medium (2) ✗ (0) Medium (4) Low (2) Medium (2) Medium (17)
24 Lopes de Souza et al. [99] 2023 Medium (3) Medium (4) Very High (3) ✗ (0) Medium (4) Low (2) ✗ (0) Medium (16)
25 Mavropoulos et al. [100] 2021 Low (2) High (5) Medium (2) ✗ (0) Medium (4) Medium (4) ✗ (0) Medium (17)
26 Mcheick et al. [101] 2016 ✗ (0) Medium (3) Low (1) Low (1) Low (1) Low (2) Low (1) Low (9)
27 Mezghani et al. [102] 2015 Low (1) Medium (3) Low (1) ✗ (0) Low (1) Low (1) Low (1) Low (8)
28 Minutolo et al. [103] 2016 Low (1) Medium (3) Very High (3) ✗ (0) ✗ (0) Low (1) Low (1) Low (9)
29 Peral et al. [104] 2018 Medium (4) Medium (3) Low (1) Medium (2) Medium (3) Low (1) ✗ (0) Medium (14)
30 Reda et al. [73] 2022 Low (1) High (5) Medium (2) Low (1) ✗ (0) Low (1) Low (1) Low (11)
31 Rhayem et al. [105] 2021 Medium (4) Very High (7) Medium (2) Low (1) Medium (3) Medium (3) Low (1) Medium (21)
32 Spoladore et al. [106] 2021 Low (2) Medium (4) Medium (2) ✗ (0) Medium (4) Low (2) ✗ (0) Medium (14)
33 Stavropoulos et al. [74] 2021 Low (1) Medium (4) Medium (2) ✗ (0) Low (2) Low (2) ✗ (0) Low (11)
34 Titi et al. [107] 2019 Medium (4) High (6) Medium (2) ✗ (0) High (5) Low (2) Low (1) Medium (20)
35 Vadillo et al. [108] 2013 Low (2) High (5) Low (1) ✗ (0) Medium (4) Low (2) ✗ (0) Medium (14)
36 Villarreal et al. [109] 2014 Low (1) Medium (4) Medium (2) ✗ (0) Medium (3) Medium (4) ✗ (0) Medium (14)
36 Villarreal et al. [109] 2014 Low (1) Medium (4) Medium (2) ✗ (0) Medium (3) Medium (4) ✗ (0) Medium (14)
37 Xu et al. [72] 2017 Low (2) Medium (3) Medium (2) ✗ (0) Low (2) Low (2) ✗ (0) Low (11)
38 Yu et al. [70] 2022 Low (2) Medium (4) Low (1) ✗ (0) High (5) Medium (4) ✗ (0) Medium (16)
39 Yu et al. [110] 2017 Low (2) Medium (4) Medium (2) Low (1) Low (1) Low (2) ✗ (0) Low (12)
40 Zafeiropoulos et al. [75] 2024 Medium (3) High (5) Very High (3) Low (1) Low (3) Low (2) ✗ (0) Medium (17)
41 Zeshan et al. [111] 2023 ✗ (0) High (5) Low (2) ✗ (0) Low (1) Low (2) ✗ (0) Low (10)
42 Zhang et al. [112] 2014 Low (2) Low (2) Medium (2) ✗ (0) Low (1) Low (2) ✗ (0) Low (9)
43 Zhou et al. [71] 2022 Medium (4) Medium (4) Low (1) Low (1) Medium (3) Low (1) Low (1) Medium (15)

### v3

@@ Appendix A Reused semantic resources
Table 17: Reused semantic resources, the systems that reuse them, and the challenges they address. Each resource includes either a hyperlink or a reference.
@@
Amigo Device Ontology [230] Provides support for the description of devices, user context, quality of service parameters, and communication protocols. [119] Interoperability; Context awareness
Association Ontology Provides basic concepts and properties for describing specific association statements to something. [131] Interoperability
Basic Formal Ontology A foundational ontology designed for use in supporting information retrieval, analysis, and integration in scientific and other domains. [116] Interoperability
Basic Geo (WGS84 - lat/long) A basic RDF vocabulary that provides a namespace for representing lat(itude), long(itude), and other information about spatially-located things. [125, 106] Interoperability; Context awareness
@@
Chronic Obstructive Pulmonary Disease Ontology (COPDology) [231] Contains concepts related to the disease, environment, equipment, patient data, and treatment. [126] Interoperability; Situation detection; Situation prediction; Decision support; Context awareness; Explainability
Context-Driven Adaptation of Mobile Services (CoDAMoS) Ontology [232] An adaptable and extensible context ontology for creating context-aware computing infrastructures by representing users, environments, hardware and software platforms, and services. [125] Interoperability; Context awareness
Core Ontology Model for Crowd-Sensing (COMCS) Ontology [233] A core ontology modelling upper-level concepts for worker selection in different crowd-sensing tasks. [125] Interoperability; Context awareness
Cyc Ontology A large knowledge base of common sense and background knowledge. [130] Interoperability
Data Analytics for Health and Connected Care (DAHCC) Ontology Captures care, patient, daily life activity recognition and lifestyle domain knowledge. [115, 107] Interoperability; Situation detection; Situation prediction; Context awareness; Explainability
@@
Diabetes Mellitus Diagnosis Ontology [234] A diabetes knowledge base that supports automatic reasoning for solving problems related to diabetes diagnosis. [116] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
@@
Diseasome [235] A.k.a. the human disease network; a graph that captures all genetic disorders and disease genes and the links between them. [105] Interoperability; Situation detection; Situation prediction; Explainability
DogOnt Ontology Offers a uniform, extensible model for smart IoT environments, including the location, capabilities, and configurations of devices, and the composition of the smart environment. [87] Interoperability; Context awareness
@@
FHIR & SSN-based T1 Diabetes Ontology (FASTO) [116] Ontology integrating FHIR standard and SSN ontology with clinical practice guidelines for real-time management of insulin for diabetes patients. [110] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
Food Ontology (FoodOn) A lightweight ontology for describing recipes, ingredients, menus, and diets. [98] Interoperability; Decision support; Explainability
Friend of a Friend (FOAF) Ontology An ontology describing persons, their activities, and their relations to other people and objects. [108, 89, 91, 93, 126, 106, 98, 100, 131] Interoperability; Context awareness
General User Model Ontology (GUMO) [236] A user model ontology for use in user-adaptive or ubiquitous computing systems. [126] Interoperability; Context awareness
@@
HealthIoT Ontology [237] Provides semantic representation of both the medical connected objects (i.e. sensors) and their data. [95] Interoperability; Context awareness
@@
International Classification of Functioning, Disability and Health (ICF) A framework for describing and organising information on functioning and disability. [91, 98] Interoperability; Situation detection; Explainability
@@
IoT-lite Ontology A lightweight ontology to represent IoT resources, entities, and services. [91, 83] Interoperability; Context awareness
IoT-Stream ontology [238] A lightweight ontology to semantically annotate stream data from IoT devices. [91] Interoperability; Context awareness
@@
Linked Data Notifications (LDN) A protocol that describes how servers (receivers) can have messages pushed to them by applications (senders), as well as how other applications (consumers) may retrieve those messages. [108] Interoperability
Linked Data Platform (LDP) Defines a set of rules for HTTP operations on web resources, some based on RDF, to provide an architecture for read-write Linked Data on the web. [108] Interoperability
Medical Subject Headings (MeSH) A vocabulary for indexing, cataloguing, and searching biomedical and health-related information. [106] Interoperability
MIMU-Wear Ontology [239] Describes wearable sensor platforms consisting of mainstream magnetic and inertial measurement units. [95] Interoperability; Context awareness
Mobile Crowd Sensing (MCS) Ontology [240] A semantic model designed to structure collected data from crowd sensing systems. [125] Interoperability; Context awareness
Mobile Object Ontology [241] Models temporal, spatial, and domain related information about mobile objects. [83] Interoperability; Context awareness
@@
Ontology of units of Measure (OM) An ontology that models concepts and relations important to scientific research, with a strong focus on units, quantities, measurements, and dimensions. [125] Interoperability; Context awareness
@@
OwlSpeak Ontology [242] Represents static and dynamic concepts related to spoken dialogue. [126] Interoperability; Context awareness
OWL-S Ontology Describes Semantic Web services for automatic service discovery, invocation, composition, and interoperation. [117, 119, 125] Interoperability; Situation detection; Situation prediction; Decision support; Context awareness
OWL-Time Ontology Describes the temporal properties of resources. [115, 125, 126, 83, 100, 131] Interoperability; Context awareness
Parkinson and Movement Disorder Ontology Captures neurological findings, treatment plans, and instruments used to evaluate various traits of Parkinson’s Disorder. [107] Interoperability; Situation detection; Situation prediction; Decision support; Context awareness; Explainability
Physical Activity Concept Ontology [243] Captures various concepts that are required to describe one’s physical activity. [95] Interoperability; Situation detection; Situation prediction; Decision support; Context awareness; Explainability
@@
Quantities and Units (QU) Ontology An ontology to model quantities and units of measurement. [91] Interoperability; Context awareness
@@
Semantic Sensor Network (SSN)/Sensor, Observation, Sample, and Actuator (SOSA) Ontology Describes sensors and their capabilities, measurement processes, observations, and deployments. [87, 114, 115, 116, 89, 91, 93, 95, 125, 83, 99, 100, 107] Interoperability; Context awareness
@@
Stream Annotation Ontology [244] An ontology for real-time semantic annotation of streaming IoT data to support dynamic integration into the Web [91] Interoperability; Context awareness
SWRL Temporal Ontology Defines defines a set of built-ins that can be used to represent temporal information in SWRL rules. [116] Interoperability; Context awareness
Systematized Nomenclature of Medicine Clinical Terms (SNOMED CT) A collection of medical terms providing codes, terms, synonyms and definitions used in clinical documentation and reporting. [87, 114, 116, 97, 123, 124, 106, 83, 100, 103] Interoperability
@@
Web Access Control Ontology Defines and enforces authorization conditions on Web resources, allowing resource owners to control who can access their data and how. [108] Interoperability
Web Annotation Ontology Structures annotations (i.e. information about a resource or associations between resources) to enable sharing and reuse across platforms. [108] Interoperability
@@ Appendix B Challenges assessment
# System Year Interoperability (score out of 6) Situation detection (score out of 3) Situation prediction (score out of 3) Decision support (score out of 6) Context awareness (score out of 7) Explainability (score out of 6) Uncertainty handling (score out of 6)
1 Akhtar et al. [109] 2022 Low (1) Medium (2) ✗ (0) Medium (3) High (5) Medium (3) Low (1)
2 Ali et al. [110] 2021 Medium (3) High (3) ✗ (0) Low (2) Medium (4) Low (2) Low (1)
3 Ali et al. [111] 2020 Low (2) Medium (2) Low (1) Medium (3) Medium (3) Low (1) Medium (2)
4 Ali et al. [112] 2018 Low (1) High (3) ✗ (0) Medium (3) Medium (3) Low (2) Medium (2)
5 Alti et al. [113] 2021 Low (1) Medium (2) Low (1) Low (1) High (5) Low (2) ✗ (0)
6 Ammar et al. [108] 2021 Medium (3) Low (1) ✗ (0) High (5) High (5) Low (2) ✗ (0)
7 Bampi et al. [87] 2025 Medium (4) Medium (2) ✗ (0) Low (2) High (5) Low (2) ✗ (0)
8 Chatterjee et al. [114] 2021 Medium (4) Medium (2) ✗ (0) Low (2) High (5) Medium (3) ✗ (0)
9 Chiang and Liang [88] 2015 Low (1) Medium (2) Low (1) Medium (4) High (5) Medium (3) Low (1)
10 De Brouwer et al. [115] 2022 Medium (4) High (3) Low (1) Medium (3) High (6) Low (2) ✗ (0)
11 El-Sappagh et al. [116] 2019 High (5) Low (1) ✗ (0) Medium (3) High (5) Low (2) ✗ (0)
12 Elhadj et al. [89] 2021 High (5) Medium (2) ✗ (0) Low (2) High (5) Medium (4) ✗ (0)
13 Esposito et al. [90] 2018 Low (1) High (3) ✗ (0) Low (2) Medium (4) Low (2) Medium (2)
14 Fenza et al. [117] 2012 Low (2) High (3) Low (1) ✗ (0) Low (2) Low (1) Medium (2)
15 Garcia-Moreno et al. [91] 2023 Medium (4) Medium (2) ✗ (0) Medium (3) High (6) Low (1) Low (1)
16 Garcia-Valverde et al. [92] 2014 ✗ (0) Low (1) ✗ (0) Low (2) Medium (3) Low (2) Low (1)
17 Hadjadj and Halimi [118] 2021 Medium (3) Low (1) ✗ (0) Medium (3) High (5) Low (2) ✗ (0)
18 Henaien et al. [93] 2020 Medium (4) Low (1) ✗ (0) Medium (3) Medium (4) Low (2) ✗ (0)
19 Hooda and Rani [94] 2020 Low (1) Medium (2) ✗ (0) Low (2) Low (2) Low (2) Low (1)
20 Hristoskova et al. [119] 2014 Medium (3) Medium (2) Low (1) Medium (3) High (6) Low (2) Low (1)
21 Hussain and Park [120] 2021 Low (2) Medium (2) ✗ (0) Low (2) Medium (3) Low (2) Low (1)
22 Ivașcu and Negru [95] 2021 Low (1) Low (1) ✗ (0) Low (1) Medium (4) Low (2) ✗ (0)
23 Ivașcu et al. [96] 2015 Low (1) Medium (2) ✗ (0) Low (1) High (5) Low (2) ✗ (0)
24 Khozouie et al. [121] 2018 Low (2) Low (1) ✗ (0) Low (1) High (5) Low (2) ✗ (0)
25 Kilintzis et al. [97] 2019 Medium (3) Low (1) ✗ (0) Low (1) Medium (4) Low (1) Low (1)
26 Kim et al. [122] 2014 Low (1) High (3) ✗ (0) Low (1) Medium (3) Low (1) ✗ (0)
27 Kordestani et al. [123] 2021 Medium (3) Medium (2) ✗ (0) Medium (4) Medium (4) Low (2) Medium (2)
28 Lopes de Souza et al. [124] 2023 Medium (3) High (3) ✗ (0) Medium (4) Medium (4) Low (2) ✗ (0)
29 Martella et al. [125] 2025 Medium (3) Low (1) ✗ (0) Low (2) High (5) Low (2) Low (1)
30 Mavropoulos et al. [126] 2021 Low (2) Medium (2) ✗ (0) Medium (4) High (5) Medium (4) ✗ (0)
31 Mcheick et al. [127] 2016 ✗ (0) Low (1) Low (1) Low (1) Medium (3) Low (2) Low (1)
32 Mezghani et al. [128] 2015 Low (1) Low (1) ✗ (0) Low (1) Medium (3) Low (1) Low (1)
33 Minutolo et al. [129] 2016 Low (1) High (3) ✗ (0) ✗ (0) Medium (3) Low (1) Low (1)
34 Peral et al. [130] 2018 Medium (4) Low (1) Medium (2) Medium (3) Medium (3) Low (1) ✗ (0)
35 Reda et al. [106] 2022 Low (1) Medium (2) Low (1) ✗ (0) High (5) Low (1) Low (1)
36 Rhayem et al. [83] 2021 Medium (4) Medium (2) Low (1) Medium (3) High (7) Medium (3) Low (1)
37 Spoladore et al. [98] 2021 Low (2) Medium (2) ✗ (0) Medium (4) Medium (4) Low (2) ✗ (0)
38 Stavropoulos et al. [99] 2021 Low (1) Medium (2) ✗ (0) Low (2) Medium (4) Low (2) ✗ (0)
39 Titi et al. [100] 2019 Medium (4) Medium (2) ✗ (0) High (5) High (6) Low (2) Low (1)
40 Vadillo et al. [101] 2013 Low (2) Low (1) ✗ (0) Medium (4) High (5) Low (2) ✗ (0)
41 Villarreal et al. [102] 2014 Low (1) Medium (2) ✗ (0) Medium (3) Medium (4) Medium (4) ✗ (0)
42 Xu et al. [105] 2017 Low (2) Medium (2) ✗ (0) Low (2) Medium (3) Low (2) ✗ (0)
43 Yu et al. [104] 2022 Low (2) Low (1) ✗ (0) High (5) Medium (4) Medium (4) ✗ (0)
44 Yu et al. [131] 2017 Medium (3) Medium (2) Low (1) Low (1) Medium (4) Low (2) ✗ (0)
45 Zafeiropoulos et al. [107] 2024 Medium (3) High (3) ✗ (0) Medium (3) High (5) Low (2) ✗ (0)
46 Zeshan et al. [132] 2023 ✗ (0) Medium (2) ✗ (0) Low (1) High (5) Low (2) ✗ (0)
47 Zhang et al. [133] 2014 Low (2) Medium (2) ✗ (0) Low (1) Low (2) Low (2) ✗ (0)
48 Zhou et al. [103] 2022 Medium (4) Low (1) Low (1) Medium (3) Medium (4) Low (1) Low (1)